A 1.30 kg squirrel runs east across the ground for 6.00 meters. She climbs 0.600 meters before running west on a limb for 3.00 meters. She jumps up .900 meters onto another branch, travels west 4.25 meters. She stops and eats a nut. Calculate net work. What is the squirrel’s gravitational potential energy?
1
Expert's answer
2020-11-12T09:17:05-0500
"W=mg(h_1+h_2)\\W=(1.3)(9.8)(0.6+0.9)=19.1\ J"